# DISCODE: Distribution-Aware Score Decoder for Robust Automatic Evaluation of Image Captioning

### v1

@@ 1 Introduction
Developing automatic evaluation metrics that closely correlate with human judgments is essential for advancing toward more human-centric artificial intelligence. For image captioning tasks, significant efforts have been devoted to designing automatic evaluation metrics, beginning with traditional methods such as BLEU (Papineni et al. 2002) and CIDEr (Vedantam, Zitnick, and Parikh 2015). Nevertheless, accurate evaluation remains challenging due to the inherent variability and subjectivity of natural language descriptions.
Recently, large vision-language models (LVLMs) have demonstrated substantial improvements in image-text alignment tasks (Liu et al. 2023a; Wang et al. 2024a; Chen et al. 2024b; Lu et al. 2024). To perform accurate numerical evaluation using LVLMs, state-of-the-art methods such as FLEUR (Lee, Park, and Kang 2024) and G-VEval (Tong et al. 2025) leverage the score smoothing technique, which generates real-valued scores by estimating the score distribution based on the output token probability distribution assigned by LVLMs. However, robustly aligning generated scores with human judgments remains difficult, especially in domain-diverse scenarios.
@@
Specifically, this paper makes two major contributions. First, we propose the Distribution-Aware Score Decoder (DISCODE), a novel test-time adaptive decoder for LVLMs that generates robust scores by leveraging a Gaussian prior distribution. Second, we introduce the Multi-domain Caption Evaluation (MCEval) benchmark, a new image captioning evaluation benchmark spanning six visual domains. In our experiments, we demonstrate that DISCODE achieves state-of-the-art performance on MCEval as well as four representative benchmarks: Flickr8k-Expert (Hodosh, Young, and Hockenmaier 2013), Flickr8k-CF (Hodosh, Young, and Hockenmaier 2013), Composite (Aditya et al. 2015), and Pascal-50S (Vedantam, Zitnick, and Parikh 2015). Our contributions are summarized as follows.
@@ 2 Related Work
Reference‑based metrics. This approach quantifies caption quality by comparing candidate captions with human-written references. Classical metrics such as BLEU (Papineni et al. 2002), ROUGE (Lin 2004), and METEOR (Banerjee and Lavie 2005) rely on ‑gram overlap. CIDEr (Vedantam, Zitnick, and Parikh 2015) incorporates TF–IDF weighting to emphasize consensus, and SPICE (Anderson et al. 2016) captures semantic structure by matching scene graphs. With advances in pretrained language models, caption similarity has become measurable in representational space rather than surface form, leading to embedding‑based metrics such as BERTScore (Zhang et al. 2020; Yi, Deng, and Hu 2020), BARTScore (Yuan, Neubig, and Liu 2021), TIGEr (Jiang et al. 2019) and ViLBERTScore (Lee et al. 2020). Finetuning-based metrics have further enhanced alignment with human judgments; examples include FAIEr (Wang et al. 2021), Polos (Wada et al. 2024), CAMScore (Cui et al. 2025), DENEB (Matsuda, Wada, and Sugiura 2024)and MLF (Gomes, Zerva, and Martins 2025).
Reference‑free metrics. To evaluate image captions without relying on human-written references, reference-free metrics leverage pre-trained vision-language models. CLIP-Score (Hessel et al. 2021) is a representative metric that measures the alignment between images and texts using CLIP embeddings (Radford et al. 2021). PAC-S (Sarto et al. 2023, 2024b) introduced positive-augmented contrastive learning, leveraging image generators to more precisely align image and text embeddings. HiFi-Score (Yao, Wang, and Chen 2024) introduced graph-based matching. Recent studies have demonstrated that LVLMs can serve as effective evaluators. To generate accurate real-valued scores, FLEUR (Lee, Park, and Kang 2024) introduced improved score smoothing for LLaVA-1.5 (Liu et al. 2023a). G-VEval (Tong et al. 2025) designed chain-of-thought prompts (Wei et al. 2022) for caption evaluation using GPT-4o. These sophisticated approaches rely on the score smoothing technique (Liu et al. 2023b), which estimates the score distribution based on the output token probability distribution. However, token probability distributions often exhibit unintended biases, such as symbolic bias, causing them to become non-unimodal. DISCODE enhances evaluation robustness by addressing this issue through minimization of the ATT loss, a novel test-time loss formulated with a Gaussian prior distribution.
@@ 3.4 Implementation Details
Figure 5 shows the default instruction prompt used in our experiments. This prompt is designed in the FLEUR framework (Lee, Park, and Kang 2024) to produce raw scores on a scale from 0.0 to 1.0. For this prompt, the target is the first decimal place of the raw score. After computing the smoothed score , the value below the decimal point in the raw score is replaced with as shown in Figure 3 (d). We build DISCODE on top of ten open-source LVLMs: LLaVA-Next-8B, -13B, -34B, -72B (Li et al. 2024), InternVL-2.5-8B, -78B (Chen et al. 2024a), Qwen2-VL-Instruct-7B, -72B (Wang et al. 2024a), CogVLM2-Chat (Hong et al. 2024) and MiniCPM-V-2.6 (Yao et al. 2024). LLaVA-Next-72B is used as a default LVLM. When using divergences other than KLD for an ablation study, we minimized the loss using the Adam optimizer for ten iterations, with an initial learning rate set to , because an analytical solution is not available.
@@ 3.5 Discussion: Why is DISCODE effective?
Previous studies (Liu et al. 2023b; Lee, Park, and Kang 2024; Tong et al. 2025) assumed that the score distribution is identical to the output token probability distribution assigned by the LVLM, i.e., . However, is not necessarily aligned with the human judgment score distribution, particularly with respect to unimodality. When a sufficiently large number of human evaluators provide scores, the resulting score distribution typically follows a unimodal distribution.111More precisely, human evaluations may become polarized into two opposing extremes, resulting in a convex but not unimodal distribution. However, such cases rarely occur within the scope of quality evaluation.In contrast, the token probability distribution may be influenced by unintended bias such as symbolic bias. In fact, the probability of the digit 0 is often overestimated as shown in Figure 4, making the distribution non-unimodal. DISCODE is effective because it addresses this issue by robustly estimating the score distribution through minimization of the ATT loss using a Gaussian prior distribution.
@@ 4 MCEval Benchmark
Reference-based |
BLEU-4 (Papineni et al. 2002) | – | 59.3 | 58.1 | 61.5 | 55.7 | 54.9 | 53.2 | 57.1 | ||||||||
| ROUGE (Lin 2004) | – | 57.9 | 56.9 | 59.5 | 54.5 | 54.4 | 50.6 | 55.6 | |||||||||
| METEOR (Banerjee and Lavie 2005) | – | 67.8 | 64.8 | 67.5 | 61.8 | 60.8 | 59.4 | 63.7 | |||||||||
@@
| BERT-S (Zhang et al. 2020) | 68.5 | 73.8 | 74.3 | 72.6 | 67.0 | 58.6 | 69.1 | ||||||||||
| Polos (Wada et al. 2024) | 81.3 | 75.0 | 77.6 | 76.8 | 74.5 | 69.0 | 75.7 | ||||||||||
Reference-free |
CLIP-S (Hessel et al. 2021) | 79.2 | 78.0 | 78.3 | 75.4 | 73.9 | 66.7 | 75.3 | |||||||||
| PAC-S (Sarto et al. 2023) | 80.7 | 71.1 | 69.7 | 67.5 | 66.8 | 58.7 | 69.1 | ||||||||||
@@
We constructed MCEval using images from DomainNet (Peng et al. 2019) and InfographicVQA (Mathew et al. 2022), covering six visual domains: real, painting, sketch, quickdraw, clipart, and infograph. Our annotation process involved three steps. First, we randomly sampled 1,000 images for each domain and generated their candidate captions using four proprietary LVLMs: GPT-4o-mini, GPT-4o, Gemini 2.0 Flash, and Claude 3.5 Sonnet. Open-source LVLMs were excluded because DISCODE depends on them. Second, we randomly selected one candidate caption per image, and human annotators revised it to form a reference caption. Third, for each image, three annotators compared two other candidate captions, selecting the better caption based on relevance, descriptiveness, correctness, and fluency. If consensus among these annotators was not reached, the image was discarded and replaced with another randomly sampled image from the same domain. Annotation tasks were completed by 81 crowdworkers via MTurk and Upwork platforms.
@@
Reference-based |
BLEU-4 (Papineni et al. 2002) | 30.6 | 30.8 | 16.9 | 8.7 | 28.3 | 30.6 | 74.0 | ||||
| ROUGE (Lin 2004) | 32.1 | 32.3 | 19.9 | 10.3 | 30.0 | 32.4 | 78.0 | |||||
| METEOR (Banerjee and Lavie 2005) | 41.5 | 41.8 | 22.2 | 11.5 | 36.0 | 38.9 | 81.1 | |||||
| CIDEr (Vedantam, Zitnick, and Parikh 2015) | 43.6 | 43.9 | 24.6 | 12.7 | 34.9 | 37.7 | 80.1 | |||||
| SPICE (Anderson et al. 2016) | 51.7 | 44.9 | 24.4 | 12.0 | 38.8 | 40.3 | 78.7 | |||||
| RefCLIP-S (Hessel et al. 2021) | 52.6 | 53.0 | 36.4 | 18.8 | 51.2 | 55.4 | 83.3 | |||||
| RefPAC-S++ (ViT-L) (Sarto et al. 2024b) | – | 57.9 | 38.8 | – | – | 61.6 | 84.7 | |||||
| Polos (Wada et al. 2024) | – | 56.4 | 37.8 | – | – | 57.6 | 86.5 | |||||
@@
Reference‑free |
UMIC (Lee et al. 2021) | – | 46.8 | – | – | – | 56.1 | 85.1 | ||||
| CLIP-S (Hessel et al. 2021) | 51.1 | 51.2 | 34.4 | 17.7 | 49.8 | 53.8 | 80.9 | |||||
| InfoMetIC+ (Hu et al. 2023) | – | 55.5 | 36.6 | – | – | 59.3 | 86.5 | |||||
| HiFi‑Score (Yao, Wang, and Chen 2024) | – | 58.4 | – | – | – | 65.8 | 83.0 | |||||
| BRIDGE (ViT‑L) (Sarto et al. 2024a) | 55.4 | 55.8 | 36.3 | 19.0 | 52.9 | 57.2 | 82.9 | |||||
| HICE‑S (Zeng et al. 2024) | 55.9 | 56.4 | 37.2 | 19.2 | 53.1 | 57.9 | 86.1 | |||||
| PAC‑S++ (ViT‑L) (Sarto et al. 2024b) | – | 57.4 | 38.5 | – | – | 62.0 | 82.4 | |||||
| CAMScore (Cui et al. 2025) | 54.8 | 55.6 | 37.5 | 19.3 | 53.4 | 57.5 | 85.8 | |||||
@@ 5.1 Experimental setting
We conducted extensive experiments to demonstrate the effectiveness of DISCODE on MCEval and four commonly used benchmarks: Flickr8k-Expert (Hodosh, Young, and Hockenmaier 2013), Flickr8k-CF (Hodosh, Young, and Hockenmaier 2013), Composite (Aditya et al. 2015) and Pascal-50S (Vedantam, Zitnick, and Parikh 2015). For MCEval, we report accuracy for each visual domain as well as the mean accuracy. For Flickr8k-Expert, Flickr8k-CF, and Composite, we utilize Kendall’s tau-b () and tau-c () as performance measures. For Pascal-50S, we report mean accuracy over the HC, HI, HM, and MM annotation types.
@@
We compare DISCODE with eight competitive baselines on MCEval, including three reference-free metrics: CLIP-S (Hessel et al. 2021), PAC-S (Sarto et al. 2023) and FLEUR (Lee, Park, and Kang 2024), and five reference-based metrics: BLEU (Papineni et al. 2002), ROUGE (Lin 2004), METEOR (Banerjee and Lavie 2005), CIDEr (Vedantam, Zitnick, and Parikh 2015), and Polos (Wada et al. 2024). FLEUR is the most relevant to our work, as it uses score smoothing.
@@ 5.3 Analysis
To investigate the compatibility of DISCODE with different LLMs, we applied it to four different LLMs within the framework of LLaVA-Next: Llama-3-8B (Dubey et al. 2024), Vicuna-13B (Chiang et al. 2023), Nous-Hermes-2-Yi-34B (Research 2024) and Qwen-1.5-72B (Team 2024). Results summarized in Table 6 indicate consistent improvements over raw scores and FLEUR in most cases, verifying the broad applicability of DISCODE. We also see that larger models exhibit higher performance, highlighting the scalability advantages.
@@
To further explore the applicability of DISCODE, we applied it to six leading LVLMs: InternVL-2.5-8B/78B (Chen et al. 2024b, a), Qwen2-VL-7B/72B-Instruct (Bai et al. 2023; Wang et al. 2024a), CogVLM2-Chat-19B (Wang et al. 2024b; Hong et al. 2024), and Mini-CPM-V-2.6 (Hu et al. 2024; Yao et al. 2024). As shown in Table 7, DISCODE consistently improves performance across these models. Among all results in Tables 6 and 7, the best performance is achieved by InternVL-2.5-78B on Flickr8k-Expert (58.1 ), by Qwen-VL-72B-Instruct on Composite (66.7 ) and MCEval (83.8%), and by LLaVA-Next-Qwen-1.5-72B on Flickr8k-CF (40.2 ) and Pascal-50S (84.5%). These results underscore that our approach is effective regardless of the pre-training methods or architectures of the LVLMs.

### v2

@@ 1 Introduction
Developing automatic evaluation metrics that closely correlate with human judgments is essential for advancing toward more human-centric artificial intelligence. For image captioning tasks, significant efforts have been devoted to designing automatic evaluation metrics, beginning with traditional methods such as BLEU (Papineni2002BLEU) and CIDEr (Vedantam2015CIDEr). Nevertheless, accurate evaluation remains challenging due to the inherent variability and subjectivity of natural language descriptions.
Recently, large vision-language models (LVLMs) have demonstrated substantial improvements in image-text alignment tasks (liu2023llava; Wang2024Qwen2VL; Chen2024InternVL; lu2024deepseekvl). To perform accurate numerical evaluation using LVLMs, state-of-the-art methods such as FLEUR (Lee2024FLEUR) and G-VEval (Tong2025GVEval) leverage the score smoothing technique, which generates real-valued scores by estimating the score distribution based on the output token probability distribution assigned by LVLMs. However, robustly aligning generated scores with human judgments remains difficult, especially in domain-diverse scenarios.
@@
Specifically, this paper makes two major contributions. First, we propose the Distribution-Aware Score Decoder (DISCODE), a novel test-time adaptive decoder for LVLMs that generates robust scores by leveraging a Gaussian prior distribution. Second, we introduce the Multi-domain Caption Evaluation (MCEval) benchmark, a new image captioning evaluation benchmark spanning six visual domains. In our experiments, we demonstrate that DISCODE achieves state-of-the-art performance on MCEval as well as four representative benchmarks: Flickr8k-Expert (Hodosh2013Flickr8kEXCF), Flickr8k-CF (Hodosh2013Flickr8kEXCF), Composite (Aditya2015Composite), and Pascal-50S (Vedantam2015CIDEr). Our contributions are summarized as follows.
@@ 2 Related Work
Reference‑based metrics. This approach quantifies caption quality by comparing candidate captions with human-written references. Classical metrics such as BLEU (Papineni2002BLEU), ROUGE (Lin2004ROUGE), and METEOR (Banerjee2005METEOR) rely on ‑gram overlap. CIDEr (Vedantam2015CIDEr) incorporates TF–IDF weighting to emphasize consensus, and SPICE (Anderson2016SPICE) captures semantic structure by matching scene graphs. With advances in pretrained language models, caption similarity has become measurable in representational space rather than surface form, leading to embedding‑based metrics such as BERTScore (Zhan2020BERTScore; Yi2020BERTScorepp), BARTScore (Weizhe2021BARTScore), TIGEr (Jiang2019TIGEr) and ViLBERTScore (Lee2020ViLBERTScore). Finetuning-based metrics have further enhanced alignment with human judgments; examples include FAIEr (wang2021faier), Polos (Wada2024PolosPolaris), CAMScore (cui2025evaluating), DENEB (Matsuda2024DENEB)and MLF (gomes-etal-2025-evaluation).
Reference‑free metrics. To evaluate image captions without relying on human-written references, reference-free metrics leverage pre-trained vision-language models. CLIP-Score (Hessel2021CLIPS) is a representative metric that measures the alignment between images and texts using CLIP embeddings (radford2021CLIP). PAC-S (Sarto2023PACS; Sarto2024PACScorepp) introduced positive-augmented contrastive learning, leveraging image generators to more precisely align image and text embeddings. HiFi-Score (yao2024hifiscore) introduced graph-based matching. Recent studies have demonstrated that LVLMs can serve as effective evaluators. To generate accurate real-valued scores, FLEUR (Lee2024FLEUR) introduced improved score smoothing for LLaVA-1.5 (liu2023llava). G-VEval (Tong2025GVEval) designed chain-of-thought prompts (wei2022chainofthought) for caption evaluation using GPT-4o. These sophisticated approaches rely on the score smoothing technique (Liu2023GEval), which estimates the score distribution based on the output token probability distribution. However, token probability distributions often exhibit unintended biases, such as symbolic bias, causing them to become non-unimodal. DISCODE enhances evaluation robustness by addressing this issue through minimization of the ATT loss, a novel test-time loss formulated with a Gaussian prior distribution.
@@ 3.4 Implementation Details
Figure 5 shows the default instruction prompt used in our experiments. This prompt is designed in the FLEUR framework (Lee2024FLEUR) to produce raw scores on a scale from 0.0 to 1.0. For this prompt, the target is the first decimal place of the raw score. After computing the smoothed score , the value below the decimal point in the raw score is replaced with as shown in Figure 3 (d). We build DISCODE on top of ten open-source LVLMs: LLaVA-Next-8B, -13B, -34B, -72B (li2024llavanext-strong), InternVL-2.5-8B, -78B (Chen2024InternVL2_5), Qwen2-VL-Instruct-7B, -72B (Wang2024Qwen2VL), CogVLM2-Chat (Hong2024CogVLM2) and MiniCPM-V-2.6 (Yao2024MiniCPMv). LLaVA-Next-72B is used as a default LVLM. When using divergences other than KLD for an ablation study, we minimized the loss using the Adam optimizer for ten iterations, with an initial learning rate set to , because an analytical solution is not available.
@@ 3.5 Discussion: Why is DISCODE effective?
Previous studies (Liu2023GEval; Lee2024FLEUR; Tong2025GVEval) assumed that the score distribution is identical to the output token probability distribution assigned by the LVLM, i.e., . However, is not necessarily aligned with the human judgment score distribution, particularly with respect to unimodality. When a sufficiently large number of human evaluators provide scores, the resulting score distribution typically follows a unimodal distribution.111More precisely, human evaluations may become polarized into two opposing extremes, resulting in a convex but not unimodal distribution. However, such cases rarely occur within the scope of quality evaluation.In contrast, the token probability distribution may be influenced by unintended bias such as symbolic bias. In fact, the probability of the digit 0 is often overestimated as shown in Figure 4, making the distribution non-unimodal. DISCODE is effective because it addresses this issue by robustly estimating the score distribution through minimization of the ATT loss using a Gaussian prior distribution.
@@ 4 MCEval Benchmark
| Reference-based | BLEU-4 (Papineni2002BLEU) | – | 59.3 | 58.1 | 61.5 | 55.7 | 54.9 | 53.2 | 57.1 | ||||||||
| ROUGE (Lin2004ROUGE) | – | 57.9 | 56.9 | 59.5 | 54.5 | 54.4 | 50.6 | 55.6 | |||||||||
| METEOR (Banerjee2005METEOR) | – | 67.8 | 64.8 | 67.5 | 61.8 | 60.8 | 59.4 | 63.7 | |||||||||
@@
| BERT-S (Zhan2020BERTScore) | 68.5 | 73.8 | 74.3 | 72.6 | 67.0 | 58.6 | 69.1 | ||||||||||
| Polos (Wada2024PolosPolaris) | 81.3 | 75.0 | 77.6 | 76.8 | 74.5 | 69.0 | 75.7 | ||||||||||
| Reference-free | CLIP-S (Hessel2021CLIPS) | 79.2 | 78.0 | 78.3 | 75.4 | 73.9 | 66.7 | 75.3 | |||||||||
| PAC-S (Sarto2023PACS) | 80.7 | 71.1 | 69.7 | 67.5 | 66.8 | 58.7 | 69.1 | ||||||||||
@@
We constructed MCEval using images from DomainNet (peng2019domainnet) and InfographicVQA (Mathew2022InfographicVQA), covering six visual domains: real, painting, sketch, quickdraw, clipart, and infograph. Our annotation process involved three steps. First, we randomly sampled 1,000 images for each domain and generated their candidate captions using four proprietary LVLMs: GPT-4o-mini, GPT-4o, Gemini 2.0 Flash, and Claude 3.5 Sonnet. Open-source LVLMs were excluded because DISCODE depends on them. Second, we randomly selected one candidate caption per image, and human annotators revised it to form a reference caption. Third, for each image, three annotators compared two other candidate captions, selecting the better caption based on relevance, descriptiveness, correctness, and fluency. If consensus among these annotators was not reached, the image was discarded and replaced with another randomly sampled image from the same domain. Annotation tasks were completed by 81 crowdworkers via MTurk and Upwork platforms.
@@
| Reference-based | BLEU-4 (Papineni2002BLEU) | 30.6 | 30.8 | 16.9 | 8.7 | 28.3 | 30.6 | 74.0 | ||||
| ROUGE (Lin2004ROUGE) | 32.1 | 32.3 | 19.9 | 10.3 | 30.0 | 32.4 | 78.0 | |||||
| METEOR (Banerjee2005METEOR) | 41.5 | 41.8 | 22.2 | 11.5 | 36.0 | 38.9 | 81.1 | |||||
| CIDEr (Vedantam2015CIDEr) | 43.6 | 43.9 | 24.6 | 12.7 | 34.9 | 37.7 | 80.1 | |||||
| SPICE (Anderson2016SPICE) | 51.7 | 44.9 | 24.4 | 12.0 | 38.8 | 40.3 | 78.7 | |||||
| RefCLIP-S (Hessel2021CLIPS) | 52.6 | 53.0 | 36.4 | 18.8 | 51.2 | 55.4 | 83.3 | |||||
| RefPAC-S++ (ViT-L) (Sarto2024PACScorepp) | – | 57.9 | 38.8 | – | – | 61.6 | 84.7 | |||||
| Polos (Wada2024PolosPolaris) | – | 56.4 | 37.8 | – | – | 57.6 | 86.5 | |||||
@@
| Reference‑free | UMIC (Lee2021UMIC) | – | 46.8 | – | – | – | 56.1 | 85.1 | ||||
| CLIP-S (Hessel2021CLIPS) | 51.1 | 51.2 | 34.4 | 17.7 | 49.8 | 53.8 | 80.9 | |||||
| InfoMetIC+ (Hu2023InfoMetIC) | – | 55.5 | 36.6 | – | – | 59.3 | 86.5 | |||||
| HiFi‑Score (yao2024hifiscore) | – | 58.4 | – | – | – | 65.8 | 83.0 | |||||
| BRIDGE (ViT‑L) (Sarto2024BRIDGE) | 55.4 | 55.8 | 36.3 | 19.0 | 52.9 | 57.2 | 82.9 | |||||
| HICE‑S (zeng2024hicescore) | 55.9 | 56.4 | 37.2 | 19.2 | 53.1 | 57.9 | 86.1 | |||||
| PAC‑S++ (ViT‑L) (Sarto2024PACScorepp) | – | 57.4 | 38.5 | – | – | 62.0 | 82.4 | |||||
| CAMScore (cui2025evaluating) | 54.8 | 55.6 | 37.5 | 19.3 | 53.4 | 57.5 | 85.8 | |||||
@@ 5.1 Experimental setting
We conducted extensive experiments to demonstrate the effectiveness of DISCODE on MCEval and four commonly used benchmarks: Flickr8k-Expert (Hodosh2013Flickr8kEXCF), Flickr8k-CF (Hodosh2013Flickr8kEXCF), Composite (Aditya2015Composite) and Pascal-50S (Vedantam2015CIDEr). For MCEval, we report accuracy for each visual domain as well as the mean accuracy. For Flickr8k-Expert, Flickr8k-CF, and Composite, we utilize Kendall’s tau-b () and tau-c () as performance measures. For Pascal-50S, we report mean accuracy over the HC, HI, HM, and MM annotation types.
@@
We compare DISCODE with eight competitive baselines on MCEval, including three reference-free metrics: CLIP-S (Hessel2021CLIPS), PAC-S (Sarto2023PACS) and FLEUR (Lee2024FLEUR), and five reference-based metrics: BLEU (Papineni2002BLEU), ROUGE (Lin2004ROUGE), METEOR (Banerjee2005METEOR), CIDEr (Vedantam2015CIDEr), and Polos (Wada2024PolosPolaris). FLEUR is the most relevant to our work, as it uses score smoothing.
@@ 5.3 Analysis
To investigate the compatibility of DISCODE with different LLMs, we applied it to four different LLMs within the framework of LLaVA-Next: Llama-3-8B (Dubey2024Llama3), Vicuna-13B (vicuna2023), Nous-Hermes-2-Yi-34B (NousResearch2024HermesYi) and Qwen-1.5-72B (QwenTeam2024qwen1.5). Results summarized in Table 6 indicate consistent improvements over raw scores and FLEUR in most cases, verifying the broad applicability of DISCODE. We also see that larger models exhibit higher performance, highlighting the scalability advantages.
@@
To further explore the applicability of DISCODE, we applied it to six leading LVLMs: InternVL-2.5-8B/78B (Chen2024InternVL; Chen2024InternVL2_5), Qwen2-VL-7B/72B-Instruct (Bai2023QwenVL; Wang2024Qwen2VL), CogVLM2-Chat-19B (Wang2023CogVLM; Hong2024CogVLM2), and Mini-CPM-V-2.6 (Hu2024MiniCPM; Yao2024MiniCPMv). As shown in Table 7, DISCODE consistently improves performance across these models. Among all results in Tables 6 and 7, the best performance is achieved by InternVL-2.5-78B on Flickr8k-Expert (58.1 ), by Qwen-VL-72B-Instruct on Composite (66.7 ) and MCEval (83.8%), and by LLaVA-Next-Qwen-1.5-72B on Flickr8k-CF (40.2 ) and Pascal-50S (84.5%). These results underscore that our approach is effective regardless of the pre-training methods or architectures of the LVLMs.